# On the complexity of estimating ground state entanglement and free energy

### v1

@@ 8 Physical Hamiltonians
For soundness suppose $\left|\psi^{\prime}\right\rangle$ satisfies $\left\langle\psi^{\prime}\right|H^{\prime}\left|\psi^{\prime}\right\rangle\leq\beta^{\prime}$. Let $\left|\phi^{\prime}\right\rangle$ be the projection of $\left|\psi^{\prime}\right\rangle$ on the low energy space of $H^{\prime}$. Note that
$\beta$ $\geq\left\langle\psi^{\prime}\right|H^{\prime}\left|\psi^{\prime}\right\rangle$ (100)
$\geq\Delta\left\langle\psi^{\prime}\right|(I-\Pi_{\leq\Delta})\left|\psi^{\prime}\right\rangle$ (101)
$\geq\Delta(1-\left\langle\psi^{\prime}\right|\Pi_{\leq\Delta}\left|\psi^{\prime}\right\rangle)$ (102)
$\left\langle\psi^{\prime}\right|\Pi_{\leq\Delta}\left|\psi^{\prime}\right\rangle$ $\geq 1-\frac{\beta}{\Delta},$ (103)
hence by the Gentle Measurement Lemma, $\|\psi^{\prime}-\phi^{\prime}\|_{1}\leq 2\sqrt{\frac{\beta}{\Delta}}$.
@@
$\left\langle\phi\right|H\left|\phi\right\rangle$ $\leq\left\langle\phi^{\prime}\right|\widetilde{V}H\widetilde{V}^{\dagger}\left|\phi^{\prime}\right\rangle$ (104)
$\leq\left\langle\phi^{\prime}\right|H^{\prime}\left|\phi^{\prime}\right\rangle+\varepsilon$ (105)
$\leq\left\langle\psi^{\prime}\right|H^{\prime}\left|\psi^{\prime}\right\rangle+2\|H\|_{\infty}\sqrt{\frac{\beta}{\Delta}}+\varepsilon.$ (106)
Hence, for an appropriate choices of $\Delta,\varepsilon$, $\left\langle\phi\right|H\left|\phi\right\rangle\leq\beta$. As we are in the NO case, this means that $S(\phi_{A})\leq t$ and as $V$ is a local isometry, $S(\chi_{A})\leq t$ for $\left|\chi\right\rangle=V\left|\phi\right\rangle$. Note that $\|\chi-\phi^{\prime}\|_{1}\leq\|V-\tilde{V}\|_{\infty}=\eta$ and that $\|\phi^{\prime}-\psi^{\prime}\|_{1}\leq 2\sqrt{\frac{\beta}{\Delta}}$, hence $\|\chi-\psi^{\prime}\|_{1}\leq\eta+2\sqrt{\frac{\beta}{\Delta}}$. By the Fannes Inequality, the entropy of $\psi^{\prime}_{A^{\prime}}$ is now close to that of $\phi_{A}$. In particular, for appropriate choices of $\eta,\Delta$, we get $S(\psi^{\prime}_{A^{\prime}})\leq t^{\prime}$.

### v2

@@ 8 Physical Hamiltonians
For soundness suppose $\rho^{\prime}$ satisfies $\operatorname{Tr}(H^{\prime}\rho)\leq\beta^{\prime}$. Let $\sigma^{\prime}=\frac{\Pi_{\leq\Delta}\rho^{\prime}\Pi_{\leq\Delta}}{\operatorname{Tr}(\Pi_{\leq\Delta}\rho^{\prime})}$ be the projection of $\rho^{\prime}$ on the low energy space of $H^{\prime}$. Note that
$\beta^{\prime}$ $\geq\operatorname{Tr}(H^{\prime}\rho^{\prime})$ (100)
$\geq\Delta\operatorname{Tr}\left((I-\Pi_{\leq\Delta})\rho^{\prime}\right)$ (101)
$\geq\Delta\left(1-\operatorname{Tr}\left(\Pi_{\leq\Delta}\rho^{\prime}\right)\right)$ (102)
$\operatorname{Tr}(\Pi_{\leq\Delta}\rho^{\prime})$ $\geq 1-\frac{\beta^{\prime}}{\Delta},$ (103)
hence by the Gentle Measurement Lemma, $\|\rho^{\prime}-\sigma^{\prime}\|_{1}\leq 2\sqrt{\frac{\beta^{\prime}}{\Delta}}$.
@@
[…]
$\leq\operatorname{Tr}(H^{\prime}\sigma^{\prime})+\varepsilon$ (105)
$\leq\operatorname{Tr}(H^{\prime}\rho^{\prime})+2\|H\|_{\infty}\sqrt{\frac{\beta^{\prime}}{\Delta}}+\varepsilon.$ (106)
Hence, for an appropriate choices of $\Delta,\varepsilon$, $\operatorname{Tr}(H\sigma)\leq\beta$. As we are in the NO case, this means that $S(\sigma_{A})\leq t$ and as $V$ is a local isometry, $S(\chi_{A})\leq t$ for $\chi=V\sigma V^{\dagger}$. Note that $\|\chi-\sigma^{\prime}\|_{1}\leq\|V-\tilde{V}\|_{\infty}=\eta$ and that $\|\sigma^{\prime}-\rho^{\prime}\|_{1}\leq 2\sqrt{\frac{\beta^{\prime}}{\Delta}}$, hence $\|\chi-\rho^{\prime}\|_{1}\leq\eta+2\sqrt{\frac{\beta}{\Delta}}$. By the Fannes Inequality, the entropy of $\rho^{\prime}_{A^{\prime}}$ is now close to that of $\chi_{A}$. In particular, for appropriate choices of $\eta,\Delta$, we get $S(\rho^{\prime}_{A^{\prime}})\leq t^{\prime}$.